# Real-Time Brain Biomechanics Prediction with Neural Operators: Toward Clinically Deployable Traumatic Brain Injury Models

### v1

@@ Abstract
Background: Traumatic brain injury (TBI) remains a major public health concern, with over 69 million cases annually worldwide. Finite element (FE) models offer high-fidelity predictions of brain deformation but are computationally expensive, requiring hours per simulation and limiting their clinical utility for rapid decision-making.
Objective: This study benchmarks state-of-the-art neural operator (NO) architectures for rapid, patient-specific prediction of brain displacement fields, aiming to enable real-time TBI modeling in clinical and translational settings.
Methods: We formulated TBI modeling as an operator learning problem, mapping subject-specific anatomical MRI, magnetic resonance elastography (MRE) stiffness maps, and demographic features to full-field 3D brain displacement predictions. Four architectures - Fourier Neural Operator (FNO), Factorized FNO (F-FNO), Multi-Grid FNO (MG-FNO), and Deep Operator Network (DeepONet) were trained and evaluated on 249 MRE datasets across physiologically relevant frequencies (20 - 90 Hz).
@@
Conclusion: NOs provide an efficient, resolution-invariant approach for predicting brain deformation, opening the door to real-time, patient-specific TBI risk assessment, clinical triage support, and optimization of protective equipment. These results highlight the potential for NO-based digital twins of the human brain, enabling scalable, on-demand biomechanical modeling in both clinical and population health contexts.
Keywords: Neural operators, Fourier Neural Operator (FNO), Traumatic brain injury, Brain biomechanics, Magnetic Resonance Elastography (MRE), Deep learning, Real-time prediction, Digital twins, Protective equipment design
@@ 1 Introduction
Traumatic brain injury (TBI) is one of the leading causes of death and disability worldwide, affecting nearly 69 million people annually and contributing to significant long-term cognitive, emotional, and physical impairments. In the United States alone, TBI accounts for over 214,000 hospitalizations and more than 200 deaths per day, with associated healthcare costs exceeding \$76 billion per year cdc_tbi_data . Importantly, TBI is not limited to acute events, up to half of patients with mild TBI or concussion experience persistent neurological deficits six months after injury. These realities underscore an urgent need for accurate, patient-specific tools that can quantify brain tissue deformation and guide clinical decision-making after head impacts matney2022scope .
Finite element (FE) models have emerged as the gold standard for characterizing brain biomechanics. When combined with high-resolution neuroimaging, anatomical T1-weighted MRI and magnetic resonance elastography (MRE) that non-invasively quantifies viscoelastic properties such as shear modulus, FE simulations can predict subject-specific brain responses under controlled loading and reproduce experimental deformation fields measured with tagged MRI hu2022advanced ; bahreinizad2023mri ; gomez2019acceleration . However, these models are computationally demanding, often requiring hours of runtime per subject due to the need to solve large, nonlinear partial differential equations over high-resolution meshes. This computational burden limits their integration into time-sensitive settings such as emergency triage, sideline concussion assessment, or rapid evaluation of helmet designs. Consequently, FE modeling remains largely a research tool rather than a clinically deployable technology griffiths2022finite .
Recent advances in machine learning (ML) have spurred efforts to accelerate TBI modeling wu2022realtime . Conventional ML approaches, such as convolutional neural networks, have been used to predict injury outcomes or scalar metrics like peak strain from imaging or kinematic features. While effective for outcome classification, these models typically lack spatial resolution and cannot produce the full-field brain deformation maps required for detailed injury risk assessment. More recent deep learning methods have trained neural networks directly on large libraries of FE simulations to predict voxelwise strain fields, achieving high accuracy but requiring extensive simulation datasets and sometimes generalizing poorly outside the training domain.
Neural operators (NOs) offer a powerful alternative. Unlike standard neural networks that approximate pointwise functions, NOs learn mappings between entire function spaces, making them resolution-invariant and able to generalize across new inputs and discretizations. Architectures such as the Deep Operator Network (DeepONet) lu2019deeponet and Fourier Neural Operator (FNO)  li2021fourier have demonstrated success in fluid dynamics, elasticity, and climate modeling, achieving orders-of-magnitude faster inference than numerical solvers while retaining high accuracy. By directly approximating the solution operator of the governing equations from labeled datasets, NOs are well-suited for modeling complex, nonlinear biomechanical responses in the brain. Despite their promise, neural operators have not been systematically evaluated for TBI biomechanics. Key questions remain regarding which operator architecture best captures the nonlinear mapping from patient-specific anatomy and material properties to full-field brain deformation, and how these models trade off between accuracy, inference speed, and generalization. Addressing these questions is critical for translating NOs from research to clinical practice.
@@
In this study, we present the first systematic benchmark of neural operator frameworks for subject-specific brain deformation prediction. Using a multimodal neuroimaging dataset of 249 magnetic resonance elastography (MRE) experiments bayly2021mr , we compare four state-of-the-art NO architectures, FNO, Factorized FNO (F-FNO), Multi-Grid FNO (MG-FNO), and DeepONet for their ability to predict three-dimensional brain displacement fields across physiologically relevant frequencies (20–90 Hz). Our goals are threefold: (1) to formulate patient-specific TBI modeling as an operator learning problem, (2) to quantify the accuracy and computational efficiency of each architecture, and (3) to identify the approaches best suited for real-time, clinically deployable TBI risk assessment and protective equipment design.
The rest of the paper is organized as follows. Section 2 describes the dataset used for the study. Section 3 provides an overview of the deep neural operator theory and the architectures used in this study. Here, we describe the formulation of FNO, F-FNO, MG-FNO and DeepONet highlighting the differences in their implementations. Section 4 presents the numerical results and comparison of the models in terms of approximation accuracy. Section 5 outlines the conclusions and future scope of our study.
@@ 2.1 Dataset Composition
The complete dataset includes multimodal imaging data from  300 participants (approximately 100 per site), encompassing a broad age range (14–80 years) and balanced sex representation. Participants were healthy volunteers recruited to capture normative biomechanical variability across age and sex. Each subject underwent one of three primary biomechanical imaging protocols, depending on site:
@@ 2.2 Data Characteristics
Representative examples of the dataset including the attributes utilized in this study are provided in Figure 1.
@@ 3 Deep Neural Operators
NOs have been successfully applied across diverse domains, including materials modeling goswami2022deep ; kag2024learning , earthquake response prediction goswami2025neural , weather forecasting kurth2023fourcastnet , fluid dynamics goswami2023physics , and biomedical imaging maier2022known . They have enabled breakthroughs in solving differential equations cao2024laplace , accelerated real-time prediction of complex dynamics kontolati2024learning ; karumuri2025physics , and coupled seamlessly with numerical solvers for hybrid approaches wang2025time . Applications also include inverse problems and parameter estimation in heterogeneous systems sarkar2025adaptive ; sarkar2025interface , time-dependent PDE modeling via physics-informed temporal operators nayak2025ti ; mandl2025physics , and resolution-independent formulations that maintain accuracy across discretizations bahmani2025resolution . Recent advances include separable physics-informed architectures that break the curse of dimensionality mandl2025separable , synergistic multitask frameworks for efficient PDE solution kumar2025synergistic , and stochastic modeling of structural systems under natural hazards goswami2025neural .
In this work, we focus on benchmarking four state-of-the-art NO architectures: FNO li2021fourier , F-FNO ffno , MG-FNO kossaifi2023multi , and DeepONet lu2019deeponet for patient-specific TBI modeling. All models were implemented in PyTorch paszke2019pytorch , and we detail their architecture, training strategies, and hyperparameter optimization in the following sections.
@@ 3.2 DeepONet Architecture
This design allows the DeepONet to efficiently learn complex mappings from multimodal brain imaging data and scalar features to full 3D displacement fields, while retaining flexibility to generalize across subjects and acquisition parameters. Architecture visualization is in Figure 4.
@@ 5 Conclusions
This study provides the first systematic benchmark of neural operator architectures for patient-specific modeling of traumatic brain injury (TBI) biomechanics. By leveraging a multimodal dataset of magnetic resonance elastography (MRE) displacement fields, we demonstrate that Fourier Neural Operator (FNO)–based models and Deep Operator Networks (DeepONet) can accurately predict full-field brain deformation across physiologically relevant frequency ranges while reducing computation time from hours to milliseconds. Among the tested architectures, the Multi-Grid FNO (MG-FNO) achieved the highest accuracy (94.3% spatial fidelity), particularly in capturing fine-scale, high-gradient displacement patterns critical for injury assessment. Factorized FNO (F-FNO) achieved the fastest training convergence, whereas DeepONet delivered the most efficient inference, suggesting potential for lightweight, real-time deployment on edge devices.
These findings represent a paradigm shift for computational neurobiomechanics. Neural operator–based surrogates bridge the gap between high-fidelity finite element models and the speed required for clinical translation, enabling real-time digital twins of the human brain. Such models could transform how TBI is managed by supporting rapid triage decisions in emergency settings, sideline concussion evaluation in sports, and personalized helmet or protective equipment design. In research and public health contexts, scalable operator-based models may facilitate population-level simulations to identify vulnerable subgroups and inform injury prevention policies.
Despite these advances, several challenges remain. The models exhibited residual spectral bias, underpredicting localized, high-frequency deformation modes most relevant to tissue-level injury risk. Future work should focus on incorporating multiscale frequency learning, hybrid physics-informed loss functions, and uncertainty quantification to improve robustness and interpretability. Extending these models to time-resolved, nonlinear impact scenarios and validating them on independent, multi-site datasets will be essential for clinical translation.
In conclusion, neural operators hold significant promise as clinically deployable tools for rapid, individualized prediction of brain biomechanics. By uniting subject-specific imaging, operator learning, and high-throughput inference, this work lays the groundwork for precision neurotrauma modeling—shifting computational biomechanics from retrospective research analyses toward proactive, real-time clinical decision support.

### v2

# Multimodal Neural Operators for Real-Time Biomechanical Modelling of Traumatic Brain Injury

@@ Abstract
Background: Traumatic brain injury (TBI) remains a major public health concern, with over 69 million cases annually worldwide. While neural operators have revolutionized scientific computing, existing architectures assume single-modality functional inputs and cannot handle the heterogeneous multimodal data required for patient-specific biomechanical modeling: anatomical imaging, scalar demographic parameters, and geometric constraints with fundamentally different statistical properties and dimensionalities.
Objective: This study introduces the first multimodal neural operator framework for biomechanics, enabling fusion of heterogeneous input modalities to predict full-field brain displacement fields for rapid, patient-specific TBI risk assessment.
Methods: We reformulated TBI modeling as a multimodal operator learning problem and proposed two fusion strategies: field projection for Fourier Neural Operator (FNO) based architectures (broadcasting scalars onto spatial grids) and branch decomposition for Deep Operator Networks (DeepONet) (separate encoding with multiplicative fusion). Four architectures (FNO, Factorized FNO (F-FNO), Multi-Grid FNO (MG-FNO), DeepONet) were extended with multimodal fusion mechanisms and evaluated on 249 in vivo Magnetic Resonance Elastography (MRE) datasets across physiologically relevant frequencies (20 to 90 Hz).
@@
Conclusion: Multimodal neural operators provide an efficient, resolution-invariant approach for fusing heterogeneous biomedical data, enabling real-time, patient-specific TBI risk assessment. Beyond TBI, this framework establishes a generalizable paradigm for operator learning in scientific domains requiring heterogeneous data fusion, including materials modeling, climate science, and precision medicine.
Keywords: Multimodal neural operators, Heterogeneous data fusion, Fourier Neural Operator (FNO), Deep Operator Network (DeepONet), Traumatic brain injury, Magnetic Resonance Elastography (MRE), Patient-specific modeling, Real-time prediction, Digital twins
@@ 1 Introduction
Neural operators have revolutionized scientific computing by learning mappings between infinite-dimensional function spaces, enabling orders-of-magnitude speedups over traditional numerical solvers for partial differential equations kovachki2023neural; azizzadenesheli2024neural. Architectures such as the Fourier Neural Operator (FNO) li2021fourier and Deep Operator Network (DeepONet) lu2019deeponet have demonstrated remarkable success across diverse domains, from fluid dynamics and weather forecasting to materials modeling and climate prediction pathak2022fourcastnet; wen2023real; goswami2022deep. However, the vast majority of neural operator research assumes single-modality functional inputs, typically scalar or vector fields defined on uniform grids such as velocity distributions or pressure fields. This assumption, while mathematically elegant and computationally tractable, fundamentally limits applicability to real-world scientific and biomedical problems where predictions depend on heterogeneous multimodal data that simultaneously incorporates high-dimensional imaging, patient-specific demographics, acquisition parameters, and geometric constraints.
Biomedical applications exemplify this multimodal challenge. Consider traumatic brain injury (TBI), affecting nearly 69 million people annually worldwide and causing over 214,000 hospitalizations and 200 deaths per day in the United States alone cdc_tbi_data. Predicting patient-specific brain tissue deformation under mechanical loading, which is critical for clinical triage and injury risk assessment, requires integrating multiple heterogeneous data sources: three-dimensional anatomical MRI capturing individual brain morphology, scalar viscoelastic tissue properties from magnetic resonance elastography (MRE), demographic features such as age and sex, acquisition parameters including loading frequency and scan direction, and binary anatomical masks constraining the solution domain matney2022scope.
Finite element (FE) models have emerged as the gold standard for characterizing brain biomechanics. When combined with high-resolution neuroimaging, FE simulations can predict subject-specific brain responses under controlled loading hu2022advanced; bahreinizad2023mri; gomez2019acceleration. However, these models are computationally demanding, often requiring hours of runtime per subject, limiting their integration into time-sensitive clinical settings griffiths2022finite. Recent machine learning approaches have attempted to accelerate TBI modeling wu2022realtime, but conventional neural networks trained on discretized FE simulations lack resolution invariance and fail to generalize across input discretizations or unseen parameter combinations.
Neural operators (NOs) offer a powerful alternative by learning mappings between entire function spaces, making them resolution-invariant and able to generalize across new inputs and discretizations. Despite their promise, neural operators have not been systematically evaluated for TBI biomechanics. More critically, almost all NO research assumes single-modality functional inputs, typically fields defined on uniform grids, yet real biomedical systems are inherently multimodal, combining both voxelwise anatomical fields and subject-specific scalar parameters.
Recent advances in conditional and parametric neural operators have begun to address heterogeneous inputs with significant limitations. Vectorized Conditional Neural Fields condition solution predictions on both initial conditions and PDE parameters hagnberger2024vectorized, while parameter-conditioned variants of U-Net and FNO have been developed for geometry-dependent problems peyvan2025fusion; oommen2024rethinking. Multi-fidelity frameworks fuse datasets of different resolution or accuracy levels howard2023multifidelity; meng2020composite. However, these approaches predominantly handle one type of conditioning, either scalar parameters appended to functional inputs or different fidelity versions of the same modality, rather than true multimodal fusion of fundamentally heterogeneous data types with distinct statistical properties, dimensionalities, and physical meanings. No prior work has systematically compared architectural strategies for multimodal operator learning or evaluated their trade-offs for biomedical prediction tasks.
@@
In this work, we introduce a novel framework for multimodal operator learning that integrates structural MRI-derived anatomy with demegraphic data to predict full-field brain deformation under oscillatory loading. Unlike prior studies that benchmark neural operators solely for PDE acceleration, we reformulate TBI biomechanics as a multimodal operator learning problem and propose a unified architecture that fuses field-based and scalar input modalities within an operator learning framework. Using a dataset of 249 MRE experiments across physiologically relevant frequencies (20-90 Hz) bayly2021mr, we compare four state-of-the-art NO architectures, FNO, Factorized FNO (F-FNO), Multi-Grid FNO (MG-FNO), and DeepONet for their ability to predict three-dimensional brain displacement fields across different frequencies (20-90 Hz). Our goals are threefold: (1) to introduce a unified operator learning framework that fuses MRI-derived anatomy with the demographgic data for biomechanical prediction, (2) to propose multimodal extensions of FNO and DeepONet that inject scalar biomechanics directly into the operator space, and (3) to identify the approaches best suited for real-time, clinically deployable TBI risk assessment to advance physics-informed neurobiomechanics.
The remainder of this paper is organized as follows: Section 2 describes the dataset. Section 3 presents the multimodal operator learning framework. Section 4 reports quantitative and qualitative evaluations. Section 5 concludes with future directions for multimodal operator learning.
@@ 2.1 Dataset Composition
The complete dataset includes multimodal imaging data from 300 participants (approximately 100 per site), encompassing a broad age range (14-80 years) and balanced sex representation. Participants were healthy volunteers recruited to capture normative biomechanical variability across age and sex. Each subject underwent one of three primary biomechanical imaging protocols, depending on site:
@@ 2.2 Data Characteristics
Representative examples of the dataset, including the attributes utilized in this study, are provided in Figure 1.
@@ 3 Neural Operators
NOs have been successfully applied across diverse domains, including materials modeling goswami2022deep; kag2024learning, earthquake response prediction goswami2025neural, weather forecasting kurth2023fourcastnet, fluid dynamics goswami2023physics, and biomedical imaging maier2022known. They have enabled breakthroughs in solving differential equations cao2024laplace, accelerated real-time prediction of complex dynamics kontolati2024learning; karumuri2025physics, and coupled seamlessly with numerical solvers for hybrid approaches wang2025time. Applications also include inverse problems and parameter estimation in heterogeneous systems sarkar2025adaptive; sarkar2025interface, time-dependent PDE modeling via physics-informed temporal operators nayak2025ti; mandl2025physics, and resolution-independent formulations that maintain accuracy across discretizations bahmani2025resolution. Recent advances include separable physics-informed architectures that break the curse of dimensionality mandl2025separable, synergistic multitask frameworks for efficient PDE solution kumar2025synergistic, and stochastic modeling of structural systems under natural hazards goswami2025neural; thiagarajan2025accelerating.
In this work, we develop multimodal extensions of four state-of-the-art neural operator architectures for patient-specific TBI modeling: FNO li2021fourier, F-FNO ffno, MG-FNO kossaifi2023multi, and DeepONet lu2019deeponet. Each architecture is augmented with multimodal fusion mechanisms to handle heterogeneous inputs combining anatomical imaging, scalar demographic parameters, and geometric constraints. All models were implemented in PyTorch paszke2019pytorch, and we detail their multimodal architectural designs, fusion strategies, training procedures, and hyperparameter optimization in the following sections.
@@ 3.2 DeepONet Architecture
The multimodal nature of our DeepONet architecture is realized through the branch network decomposition strategy. Each input modality, including anatomical imaging via CNN branch and scalar features (scan direction, vibration frequency, sex, brain volume, age) via individual FNN branches, is independently encoded into a 300-dimensional latent representation by its dedicated subnetwork. These modality embeddings are fused via elementwise multiplication to create a unified multimodal representation capturing the joint influence of all inputs on brain displacement. This multiplicative fusion allows the network to learn nonlinear interactions between heterogeneous data types, enabling each modality to modulate others. The resulting combined branch representation encodes spatial anatomical features and demographic acquisition parameters, ensuring predictions are conditioned on complete multimodal subject context. Architecture visualization is in Figure 4.
@@ 5 Conclusions
This work introduces the first multimodal neural operator framework for patient-specific modeling of traumatic brain injury biomechanics. The central innovation lies in addressing the fundamental challenge of heterogeneous data fusion: integrating three-dimensional anatomical imaging, scalar demographic and acquisition parameters, and geometric constraints within a unified operator learning framework. Unlike prior surrogate modeling efforts relying on single-input field representations, our multimodal formulation enables neural operators to jointly learn from fundamentally different data types with distinct statistical properties, dimensionalities, and physical meanings. This represents a critical advancement for biomechanical modeling where predictions inherently depend on interactions between anatomical morphology, tissue properties, and individual patient characteristics.
We proposed and systematically compared two architectural strategies for multimodal fusion: field projection for FNO-based architectures and branch decomposition for DeepONet-based architectures. Through evaluation of four operator architectures on 249 in vivo MRE experiments, we demonstrated that multimodal neural operators accurately predict full-field 3D brain displacement while achieving inference times three to six orders of magnitude faster than conventional finite element solvers. Multi-Grid FNO achieved the highest accuracy with 94.3% spatial fidelity, particularly excelling in capturing fine scale, high gradient displacement patterns critical for injury assessment. Factorized FNO achieved the fastest training convergence, while DeepONet delivered the most efficient inference suitable for edge deployment. These results establish that multimodal fusion strategy choice significantly impacts both accuracy and computational efficiency.
These findings represent a paradigm shift for computational neurobiomechanics. Multimodal neural operator based surrogates bridge the gap between high-fidelity finite element models and clinical translation speed, enabling real time digital twins of the human brain. Beyond clinical applications in emergency TBI triage and personalized protective equipment design, this framework establishes a generalizable paradigm for scientific domains requiring heterogeneous data fusion, including materials modeling, climate science, and precision medicine.
Despite these advances, challenges remain. Models exhibited residual spectral bias, underpredicting localized, high frequency deformation modes most relevant to tissue-level injury risk. Future work should incorporate multiscale frequency learning, hybrid physics informed loss functions, and uncertainty quantification. Investigating attention-based fusion mechanisms may further enhance multimodal integration. Extending models to time resolved, nonlinear impact scenarios and validating on independent, multisite datasets will be essential for clinical translation.
@@ 6 Acknowledgments
Ethics Statement
This study analyzed publicly available, de-identified data from the Brain Biomechanics Imaging Resources (BBIR) repository (https://www.nitrc.org/projects/bbir). All original data were collected under IRB-approved protocols at Washington University in St. Louis, the University of Delaware, and the National Institutes of Health, with written informed consent obtained from all participants (parental consent for minors). Data were de-identified in accordance with HIPAA regulations. As this constituted secondary analysis of existing, publicly available, anonymized data, no additional IRB approval was required at Johns Hopkins University. All procedures complied with relevant institutional guidelines and the GNU GPL v3.0 license. Original ethics approvals are documented in Bayly et al. (2021).